# Splines-Based Feature Importance in Kolmogorov-Arnold Networks: A Framework for Supervised Tabular Data Dimensionality Reduction

### v1

@@ Abstract
High-dimensional datasets require effective feature selection to improve predictive performance, interpretability, and robustness. We propose and evaluate feature selection methods for tabular datasets based on Kolmogorov-Arnold networks (KANs), which parameterize feature transformations through splines, enabling direct access to interpretable importance measures. We introduce four KAN-based selectors (KAN-L1, KAN-L2, KAN-SI, KAN-KO) and compare them against classical baselines (LASSO, Random Forest, Mutual Information, SVM-RFE) across multiple classification and regression tabular dataset benchmarks. Average (over three retention levels: 20%, 40%, and 60%) F1 scores and $R^{2}$ score results reveal that KAN-based selectors, particularly KAN-L2, KAN-L1, KAN-SI, and KAN-KO, are competitive with and sometimes superior to classical baselines in structured and synthetic datasets. However, KAN-L1 is often too aggressive in regression, removing useful features, while KAN-L2 underperforms in classification, where simple coefficient shrinkage misses complex feature interactions. KAN-L2 and KAN-SI provide robust performance on noisy regression datasets and heterogeneous datasets, aligning closely with ensemble predictors. In classification tasks, KAN selectors such as KAN-L1, KAN-KO, and KAN-SI sometimes surpass the other selectors by eliminating redundancy, particularly in high-dimensional multi-class data. Overall, our findings demonstrate that KAN-based feature selection provides a powerful and interpretable alternative to traditional methods, capable of uncovering nonlinear and multivariate feature relevance beyond sparsity or impurity-based measures.
Keywords: Supervised learning, Tabular data, Feature selection, Kolmogorov-Arnold Networks
@@ 1 Introduction
High-dimensional datasets pose challenges to both interpretability and computational efficiency in machine learning. Feature selection addresses these issues by identifying and retaining the most informative features. Traditional supervised approaches to tabular data feature selection fall into three broad categories: filters, wrappers, and embedded methods. Filter methods include statistical criteria such as Pearson correlation and mutual information (MI) (Peng et al., 2005), which are fast and model-agnostic but ignore feature interactions and often select redundant features. Embedded methods integrate selection into the learning algorithm itself, such as Least Absolute Shrinkage and Selection Operator(LASSO) (Tibshirani, 1996a), spline-penalized regression (Marx and Eilers, 1996), tree-based models like Random Forests (Breiman, 2001), and gradient-boosted methods such as XGBoost (Chen and Guestrin, 2016), all of which provide model-aware importance scores but may sacrifice interpretability in complex non-linear settings. Wrapper methods, such as Recursive Feature Elimination (RFE) for Support Vector Machines (SVMs) (Guyon et al., 2002), directly search for subsets of features by iteratively training and evaluating models; these often yield strong predictive performance but are computationally expensive and prone to overfitting. Recent heuristic ensemble strategies such as the Minimum Union (Min) method (Radovic et al., 2017) have also been proposed to combine outputs of multiple selectors, improving stability but adding complexity. Finally, unsupervised projection techniques like Principal Component Analysis (PCA) (Jolliffe, 2002) reduce dimensionality by constructing orthogonal linear combinations of the original variables, which improves computational efficiency but discards direct interpretability and neglects target information. Overall, while traditional methods offer a rich toolbox for feature selection, they involve trade-offs among computational cost, predictive performance, stability, and interpretability. Filter methods are fast but myopic, wrappers are accurate but expensive, and embedded methods are model-aware but often opaque. These limitations motivate the exploration of new architectures, such as Kolmogorov-Arnold Networks (KANs), which offer both model-awareness and direct interpretability by parameterizing each feature transformation as a trainable spline.
Kolmogorov-Arnold Networks (KANs) are a recently proposed neural architecture (Liu et al., 2024b) inspired by the Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function can be expressed as a finite superposition of univariate functions and addition. KANs operationalize this by replacing traditional linear weights with trainable spline functions. Each input feature is mapped through edge-specific splines, aggregated by hidden nodes, and then transformed by learnable outer splines. This “weights-as-functions” paradigm offers both strong approximation power with compact models and intrinsic interpretability: irrelevant features tend to learn nearly flat splines, while informative features produce strong nonlinear mappings. As such, KANs naturally provide mechanisms for assessing feature relevance, making them attractive for supervised feature selection in tabular data. Despite its promise, the use of KANs for feature selection remains largely unexplored. The most explicit attempt is that of Zheng et al. (2025), who proposed a KAN-assisted Whale Optimization Algorithm (KAN-WOA) for high-dimensional medical data, where KAN acted as a surrogate fitness evaluator in evolutionary search. While effective in identifying compact biomarker subsets, this approach inherits the computational cost and sensitivity of meta-heuristic optimization and leverages KAN only as a surrogate, leaving its interpretability underutilized. Other preliminary studies point to similar potential: Wang et al. (2025) used spline visualizations to manually prune redundant features without loss of accuracy, and Liu et al. (2024a) employed a two-layer KAN as a feature extractor for sensor data, outperforming CNN-based alternatives with fewer parameters.
Taken together, these works remain proof-of-concept, underscoring both the promising potential of KAN for supervised feature selection, but also highlighting important gaps: the absence of automated, principled methods for ranking or pruning features directly within KAN, and a lack of large-scale evaluations on diverse tabular datasets. Addressing these gaps is critical to establish KAN as a practical tool for supervised feature selection. consequently, we present the first systematic investigation of Kolmogorov-Arnold Networks (KANs) structures as supervised learning feature selectors for tabular data, providing importance score-free leakage via cross validation. We develop and evaluate four complementary KAN-based importance measures: (i) coefficient-based norms of spline weights (KAN-L1 and KAN-L2), which capture the magnitude of each feature’s learned transformations; (ii) a sensitivity integral approach (KAN-SI), which quantifies the expected gradient of the output with respect to each input; and (iii) a knock-out strategy (KAN-KO), which measures the increase in predictive loss when a feature’s contribution is ablated. Last but not least, we benchmark these four KAN-based feature selectors against six state-of-the-art filter, wrapper, and embedded techniques across eight real-world datasets, assessing predictive performance, selection stability, interpretability, and robustness to irrelevant variables.
@@
The paper is organized as follows. Section 2 surveys related work on feature selection methods. Section 3 provides background and introduces Kolmogorov-Arnold Networks (KANs). Section 4 presents the methodology: coefficient-based feature importance (KAN-L1, KAN-L2), KAN-Knockout (KAN-KO), and KAN-Sensitivity Integral (KAN-SI), and it specifies the assessment pipeline from feature selection to prediction (predictors, evaluation metrics, and leakage-safe cross-validation) together with the baseline methods. Section 5 describes the dataset, experimental results and analysis along with a discussion of different results.
@@ 2 Related Work
One of the most widely used approaches is the Least Absolute Shrinkage and Selection Operator (LASSO) (Tibshirani, 1996b), which adds an $L^{1}$ penalty to the regression objective to shrink unimportant coefficients to zero. LASSO produces sparse, interpretable models and is computationally efficient, making it attractive for high-dimensional tabular problems. However, when predictors are highly correlated, it tends to retain only one from the group, potentially discarding other relevant variables (Zou and Hastie, 2005). To address LASSO’s inability to model complex nonlinear relationships while enabling flexible yet interpretable modeling of nonlinear effects, Generalized Additive Models (GAMs) (Wood and Augustin, 2002) were proposed for variable selection by fitting penalized splines to the covariates. However, this process is computationally demanding especially when extended to very high-dimensional feature spaces. Tree-based ensemble methods, particularly Random Forests (Breiman, 2001), provide feature importance scores based on mean decrease in impurity or accuracy and can naturally capture feature interactions. While more robust to multicolinearity than linear models, their importance measures can be biased toward variables with more categories or greater variance, and the rankings may not always transfer well to other model types (Strobl et al., 2007). Wrapper approaches such as Recursive Feature Elimination (RFE) (Guyon et al., 2002) attempt to overcome such biases by iteratively training a model, removing the least important features, and repeating until an optimal subset remains. This approach can yield strong predictive accuracy and adapt to various learners, but its repeated training cycles make it computationally expensive for large tabular datasets or complex models. Information-theoretic methods like Mutual Information (MI) ranking (Peng et al., 2005) take a model-agnostic route, measuring the dependency between each feature and the target variable. MI can detect nonlinear dependencies missed by purely linear methods and avoids repeated model training. However, because it evaluates features individually, it ignores redundancy, often selecting correlated variables that add little marginal information.
Introduced by Chen and Guestrin (2016), Extreme Gradient Boosting (XGBoost) rapidly became one of the most widely used tree-ensemble methods, valued for its computational efficiency and built-in feature importance metrics. Tree-based models like XGBoost naturally perform embedded feature selection by evaluating candidate splits during training, and they produce importance scores (e.g., information gain or split frequency) that can be used to rank and filter features. Over the past decade, this capability has been extensively leveraged. For instance, Kasongo and Sun (2020) applied XGBoost to a network intrusion dataset, reducing the dimensionality from 42 to 19 features based on importance scores and achieving substantial gains in classification accuracy. Building on this, Lundberg and Lee (2017) introduced SHAP (SHapley Additive exPlanations), a unified and theoretically grounded framework for quantifying each feature’s contribution to a model’s predictions. SHAP-based feature selections typically rank variables by the magnitude of their contribution to the prediction, providing greater interpretability than native importance scores. These approaches have been adopted in diverse applications for example, in medical settings to identify clinically relevant predictors while also explaining model decisions. Empirical comparisons, however, suggest trade-offs: Wang et al. (2024) found that simple importance-based selection using XGBoost’s native scores often outperformed SHAP-based selection in terms of both AUPRC and efficiency on credit fraud data, given that SHAP computation incurs additional overhead.
@@ 3.1 Kolmogorov-Arnold Networks
Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024b; Akazan et al., 2025) are inspired by Kolmogorov’s superposition theorem, which states that any multivariate continuous function $f(x_{1},\dots,x_{n})$ can be represented as a finite composition of continuous univariate functions. In the KAN formulation, a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is modeled as
@@
where $\psi_{ij}:[0,1]\to\mathbb{R}$ are univariate feature transformations (implemented via spline expansions), and $\phi_{j}:\mathbb{R}\to\mathbb{R}$ are continuous nonlinearities. Each input feature $x_{i}$ is expanded in a spline basis as
@@
where $\{B_{k}\}_{k=1}^{K}$ are fixed basis functions (e.g., B-splines), and $w_{ijk}$ are trainable coefficients. These parameters are learned end-to-end by backpropagation.
@@
Let $x\in\mathbb{R}^{d}$ be an input vector and $m$ the number of outputs. In this case, a KAN layer computes
@@
where
  • $W_{\text{base}}\in\mathbb{R}^{m\times d}$ is the base weight matrix,
  • $\phi:\mathbb{R}\to\mathbb{R}$ is an activation applied component-wise to $x$,
  • $b_{j}(x_{j})\in\mathbb{R}^{K}$ is the B-spline basis expansion of $x_{j}$,
  • $W_{\text{spline}}^{(j)}\in\mathbb{R}^{m\times K}$ are the spline weights for feature $j$.
Stacking all spline bases into one long vector
$b(x)\;=\;\big[\,b_{1}(x_{1})^{\top}\;\;b_{2}(x_{2})^{\top}\;\;\cdots\;\;b_{d}(x_{d})^{\top}\big]^{\top}\in\mathbb{R}^{dK}$ (4)
and flattening spline weights into a single matrix
$W_{\text{spline}}\;=\;\big[\,W_{\text{spline}}^{(1)}\;\;W_{\text{spline}}^{(2)}\;\;\cdots\;\;W_{\text{spline}}^{(d)}\,\big]\;\in\;\mathbb{R}^{m\times dK},$ (5)
the layer reduces to
@@
For $X\in\mathbb{R}^{n\times d}$ (rows are samples), define $\Phi(X)\in\mathbb{R}^{n\times d}$ with $(\Phi(X))_{ij}=\phi(X_{ij})$. For the spline part, each feature column $X_{j}$ is expanded into its $K$-dimensional basis $b_{j}(X_{j})\in\mathbb{R}^{n\times K}$. Concatenating these expansions across all features gives
$B(X)\;=\;\big[\,b_{1}(X_{1})\;\;b_{2}(X_{2})\;\;\cdots\;\;b_{d}(X_{d})\,\big]\;\in\;\mathbb{R}^{n\times dK}.$ (7)
Then the outputs $Y\in\mathbb{R}^{n\times m}$ are

### v2

@@ Abstract
Feature selection is a key step in many tabular prediction problems, where multiple candidate variables may be redundant, noisy, or weakly informative. We investigate feature selection based on Kolmogorov-Arnold Networks (KANs), which parameterize feature transformations with splines and expose per-feature importance scores in a natural way. From this idea we derive four KAN-based selection criteria (coefficient norms, gradient-based saliency, and knockout scores) and compare them with standard methods such as LASSO, Random Forest feature importance, Mutual Information, and SVM-RFE on a suite of real and synthetic classification and regression datasets. Using average F1 and $R^{2}$ scores across three feature-retention levels (20%, 40%, 60%), we find that KAN-based selectors are generally competitive with, and sometimes superior to, classical baselines. In classification, KAN criteria often match or exceed existing methods on multi-class tasks by removing redundant features and capturing nonlinear interactions. In regression, KAN-based scores provide robust performance on noisy and heterogeneous datasets, closely tracking strong ensemble predictors; we also observe characteristic failure modes, such as overly aggressive pruning with an $\ell_{1}$ criterion. Stability and redundancy analyses further show that KAN-based selectors yield reproducible feature subsets across folds while avoiding unnecessary correlation inflation, ensuring reliable and non-redundant variable selection. Overall, our findings demonstrate that KAN-based feature selection provides a powerful and interpretable alternative to traditional methods, capable of uncovering nonlinear and multivariate feature relevance beyond sparsity or impurity-based measures. [^1] The code can be found at https://github.com/AngeClementAkazan/Feature-Selection-Using-KAN
@@ 1 Introduction
Feature selection plays a central role in many machine learning pipelines, where models must cope with redundant, noisy, or weakly informative input variables. By identifying and retaining only the most relevant features, it can improve both interpretability and computational efficiency. Classical approaches to feature selection are often grouped into three categories: filter methods, which rank features using data-driven scores independent of the predictor; wrapper methods, which evaluate subsets of features through repeated model training; and embedded methods, which perform selection as part of the model fitting procedure itself.
@@
Filter methods include statistical criteria such as Pearson correlation and mutual information (MI) (Peng et al., 2005), which are fast and model-agnostic but ignore feature interactions and often select redundant features. Embedded methods integrate selection into the learning algorithm itself, such as Least Absolute Shrinkage and Selection Operator(LASSO) (Tibshirani, 1996a), spline-penalized regression (Marx and Eilers, 1996), tree-based models like Random Forests (Breiman, 2001), and gradient-boosted methods such as XGBoost (Chen and Guestrin, 2016), all of which provide model-aware importance scores but may sacrifice interpretability in complex non-linear settings. Wrapper methods, such as Recursive Feature Elimination (RFE) for Support Vector Machines (SVMs) (Guyon et al., 2002), directly search for subsets of features by iteratively training and evaluating models; these often yield strong predictive performance but are computationally expensive and prone to overfitting. Recent heuristic ensemble strategies such as the Minimum Union (Min) method (Radovic et al., 2017) have also been proposed to combine outputs of multiple selectors, improving stability but adding complexity. Finally, unsupervised projection techniques like Principal Component Analysis (PCA) (Jolliffe, 2002) reduce dimensionality by constructing orthogonal linear combinations of the original variables, which improves computational efficiency but discards direct interpretability and neglects target information. Overall, while traditional methods offer a rich toolbox for feature selection, they involve trade-offs among computational cost, predictive performance, stability, and interpretability. Filter methods are fast but myopic, wrappers are accurate but expensive, and embedded methods are model-aware but often opaque. These limitations motivate the exploration of new architectures, such as Kolmogorov-Arnold Networks (KANs), which offer both model-awareness and direct interpretability by parameterizing each feature transformation as a trainable spline.
Kolmogorov-Arnold Networks (KANs) are a recently proposed neural architecture (Liu et al., 2024b) inspired by the Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function can be expressed as a finite superposition of univariate functions and addition. KANs operationalize this idea by replacing traditional scalar weights with trainable spline functions: each input feature is passed through one-dimensional, edge-specific splines, aggregated at hidden nodes, and then transformed by learnable outer splines. This “weights-as-functions” paradigm yields compact yet expressive models and, crucially for feature selection, a structured parameterization in which all parameters associated with a given feature form an explicit block. Intuitively, if the learned splines attached to feature $j$ are nearly flat, then the network’s output varies little with respect to that coordinate, whereas large or highly curved splines signal a strong, nonlinear dependence. KANs therefore provide a natural basis for defining feature-importance scores directly from a trained model.
Despite this appealing structure, existing work on KANs for feature selection remains sparse and does not fully exploit these properties. Zheng et al. (2025) use KAN as a surrogate fitness evaluator within a Whale Optimization Algorithm for high-dimensional medical data, inheriting the computational cost and sensitivity of meta-heuristic search and treating KAN largely as a black-box scorer. Other preliminary studies rely on manual spline inspection to prune features (Wang et al., 2025) or use KAN as a front-end encoder for IMU time series(Liu et al., 2024a), without turning its spline parameters and gradients into systematic importance measures.
In this work, we close this gap by introducing a family of KAN-based feature-importance criteria that explicitly leverage the weights-as-functions view. We derive coefficient-based norms of spline parameters (KAN-L1, KAN-L2) to capture the global magnitude of each feature’s learned univariate transformation, a sensitivity integral based on input gradients (KAN-SI) to quantify local influence, and a knock-out score (KAN-KO) that measures the increase in loss when a feature’s spline block is ablated. We then systematically evaluate these KAN-based selectors as supervised feature-selection methods on tabular classification and regression benchmarks, comparing them with representative filter, wrapper, and embedded baselines. Our analysis covers predictive performance (macro-$F_{1}$ and $R^{2}$), robustness to redundancy (average pairwise correlation among selected features), stability (Jaccard similarity of selected sets across the five cross-validation folds), and interpretability (plots of KAN layer responses and class logits for top-ranked features)  [^2] See appendix Appendix. Empirically, KAN-L2, KAN-SI, and KAN-KO are competitive with or superior to classical baselines on structured and synthetic datasets, while remaining robust on noisy real-world tasks; KAN-L1 can be highly effective in some classification settings but tends to over-prune in regression. Overall, our results indicate that KAN-based selectors provide a practical and interpretable alternative to traditional methods.
@@
The paper is organized as follows. Section 2 surveys the related work on feature selection methods. Section 3 provides background and introduces Kolmogorov-Arnold Networks (KANs). Section 4 presents the methodology: coefficient-based feature importance (KAN-L1, KAN-L2), KAN-Knockout (KAN-KO), and KAN-Sensitivity Integral (KAN-SI), and it specifies the assessment pipeline from feature selection to prediction (predictors, evaluation metrics, and leakage-safe cross-validation) together with the baseline methods. Section 5 describes the dataset, experimental results and analysis along with a discussion of different results.
@@ 2 Related Work
One widely used embedded method is the Least Absolute Shrinkage and Selection Operator (LASSO) (Tibshirani, 1996b), which adds an $\ell_{1}$ penalty to shrink small coefficients to zero, yielding sparse and interpretable linear models at modest computational cost. However, in the presence of strong collinearity it typically keeps only one predictor from a correlated group, potentially discarding relevant variables (Zou and Hastie, 2005). Generalized Additive Models (GAMs) (Wood and Augustin, 2002) extend linear models with penalized splines to capture nonlinear effects while retaining some interpretability, but fitting and selecting splines becomes computationally demanding as dimensionality grows. Tree-based ensembles such as Random Forests (Breiman, 2001) provide impurity-based importance scores and naturally capture interactions, yet their importance can be biased toward high-variance or high-cardinality features and need not transfer well to other model classes (Strobl et al., 2007). Wrapper methods like Recursive Feature Elimination (RFE) (Guyon et al., 2002) iteratively retrain a model while pruning low-importance features, often achieving strong accuracy but at significant computational cost. Finally, information-theoretic filter methods such as Mutual Information (MI) ranking (Peng et al., 2005) can detect nonlinear dependencies without repeated model training, but because they score features individually, they do not account for redundancy and tend to select correlated variables with overlapping information.
Extreme Gradient Boosting (XGBoost) (Chen and Guestrin, 2016) is a widely used tree-ensemble method noted for its efficiency and built-in feature importance metrics. During training, XGBoost performs embedded feature selection by evaluating candidate splits and produces importance scores (e.g., gain or split counts) that can be used to rank and filter variables; this has been exploited in numerous applications, such as reducing a 42-dimensional intrusion-detection dataset to 19 informative features with improved accuracy (Kasongo and Sun, 2020). SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017) extends this idea by providing theoretically grounded, instance-level attributions whose magnitudes can be aggregated for feature ranking, often yielding more interpretable importance profiles and supporting domain-facing explanations (e.g., in medical risk models). Empirical studies, however, highlight trade-offs: Wang et al. (2024) report that simple XGBoost importance-based selection can outperform SHAP-based selection in both AUPRC and computational efficiency on credit fraud data, reflecting the additional overhead of SHAP computation.
@@
Beyond the original KAN formulation(Liu et al., 2024b; Akazan et al., 2025), recent variants such as GKAN lifts KANs to graph-structured domains, learning node/graph functions via Kolmogorov-Arnold compositions on topology-aware inputs (Kiamari et al., 2024). In biomedicine, interpretable Graph KANs fuse multi-omics with graph priors to achieve multi-cancer classification and biomarker discovery, emphasizing model transparency through graph-aware architectures (Alharbi et al., 2025). Complementarily, KAN-guided Whale Optimization uses KAN outputs to steer a metaheuristic for feature selection on medical datasets, reporting empirical gains but relying on heuristic search rather than first-principles criteria (Zheng et al., 2025). Our contribution is orthogonal: we introduce a generic KAN feature-importance framework (KAN-SI,KO, L1 and L2) that derives importance from explicit chain-rule sensitivities and spline energy, treats categoricals in a reference-invariant manner, and separates relevance from redundancy through a derivative-Gram audit enabling diversity-aware selection.
@@ 3 Background
This part provides a detailed analysis of the vanilla KAN, which we used for our study.
@@ 3.1 Kolmogorov-Arnold Networks
KANs (Liu et al., 2024b; Akazan et al., 2025) are inspired by Kolmogorov’s superposition theorem, which states that any multivariate continuous function $f(x_{1},\dots,x_{n})$ can be represented as a finite composition of continuous univariate functions. KAN models $f:\mathbb{R}^{n}\to\mathbb{R}$ as:
@@
where $\psi_{i,j}:[0,1]\to\mathbb{R}$ are the univariate functions (B-splines in this case) $\phi_{j}:\mathbb{R}\to\mathbb{R}$ are continuous functions.
Each input feature $x_{i}$ is passed through a spline basis transformation:
@@
where $B_{k}$} is a fixed spline basis (e.g., B-splines or sinusoidal), and $w_{ijk}$ are learnable coefficients. These coefficients are learned via backpropagation.
@@
Let $x\in\mathbb{R}^{d}$ be an input vector and $m$ the number of outputs. A KAN layer produces
@@
where: $W_{\text{base}}\in\mathbb{R}^{m\times d}$ is the base (linear) weight matrix; $\phi:\mathbb{R}^{d}\!\to\!\mathbb{R}^{d}$ is applied componentwise (e.g. $\mathrm{SiLU}$), so $(\phi(x))_{j}=\phi(x_{j})$; $b_{j}(x_{j})\in\mathbb{R}^{K}$ is the $K$-dimensional B-spline basis vector for feature $j$; $W_{\text{spline}}^{(j)}\in\mathbb{R}^{m\times K}$ are the spline weights attached to feature $j$.
Stacking all spline bases into a single vector
$b(x)=\big[\,b_{1}(x_{1})^{\top}b_{2}(x_{2})^{\top}\cdots b_{d}(x_{d})^{\top}\big]^{\top}\in\mathbb{R}^{d\times K},$ (4)
and concatenating weights
$W_{\text{spline}}=\big[\,W_{\text{spline}}^{(1)}W_{\text{spline}}^{(2)}\cdots W_{\text{spline}}^{(d)}\big]\in\mathbb{R}^{m\times d\times K},$ (5)
the layer is simply
@@
For $X\in\mathbb{R}^{n\times d}$ (rows are samples), define $\Phi(X)\in\mathbb{R}^{n\times d}$ by $(\Phi(X))_{ij}=\phi(X_{ij})$, and $B(X)\in\mathbb{R}^{n\times d\times K}$ by concatenating $b_{j}(X_{:j})$ columnwise. Then the output matrix $Y\in\mathbb{R}^{n\times m}$ is